# PAT: Accelerating LLM Decoding via Prefix-Aware Attention with Resource Efficient Multi-Tile Kernel

### v2

@@ 8.6. Overhead Analysis
Figure 14. Overhead of pack scheduler and preprocess latency of serving system.:
The primary overhead introduced by PAT comes from the pack scheduler, which packs decode batches into CTAs at runtime. To mitigate its overhead, as described in §5.1, we design a lazy update mechanism that reduces the schedule’s triggering frequency without affecting correctness and overlaps its execution with the serving system’s preprocessing stage. To evaluate the overhead, we measure both the scheduling latency and the preprocessing latency under the toolagent and conversation traces at a request rate of 8 and 5 req/s. As shown in Figure 14, the average scheduling latency is lower than the preprocessing latency by 81.6% and 88.8%, respectively. Consistently, the P99 TPOT in Figure 11 c1- c4 shows 19.4–93.4% reduction compared with the baselines. Therefore, when running on an asynchronous CPU thread, the pack scheduler will not introduce additional end-to-end latency, demonstrating the effectiveness of the lazy update mechanism in PAT.
@@ 9. Discussion
Prospects and Limitations. PAT leverages cross-request prefix sharing to improve bandwidth utilization and cut redundant global memory accesses in memory-bound decode attention. Its effectiveness depends on three factors: (1) hardware compute-to-bandwidth ratio. As GPUs become more compute-dominant (e.g., NVIDIA V100 to B200: 139 to 312 FLOP/Byte), the gap between compute and memory widens, so memory-focused designs like PAT become increasingly valuable; (2) model architecture. PAT yields large gains for common architectures that retain a KV cache (MHA, GQA), but benefits may shrink for architectures that compress or remove KV state (MLA (liu2024deepseek), linear attention (shen2021efficient), MLKV (zuhri2024mlkv)); and (3) prefix-sharing ratio in the batch. High concurrency with cross-request shared prefixes amplifies PAT ’s advantage, whereas small batches or workloads without shared prefixes limit the improvement (see Figure 10 1, 1920).
@@ 10. Related Work
KV Cache Related Optimization. Prior work reduced KV cache memory and fragmentation: FasterTransformer (nvidia2025fastertransformer) provided an early static-batching implementation, Orca (yu2022orca) improved utilization with continuous batching but relied on pre-allocated caches, which leads to memory fragmentation. vLLM introduced a paged KV cache to cut waste by virtual memory management (kwon2023pagedattention). SGLang’s prefix-reuse lets shared prefix KV blocks be reused across requests and is widely used in production (zheng2024sglang). Recent systems (qin2024mooncake; lee2024infinigen) further extend KV capacity by offloading KV blocks between GPU and CPU or NVMe storage, which reduces on-GPU memory pressure at the cost of additional data movement. These approaches shrink KV memory costs but do not accelerate decode attention, which dominates latency as context and output lengths grow. PAT is orthogonal to these approaches because its pack scheduler only relies on logical KV block IDs, and the serving system transfers the required KV blocks to GPU memory before the attention kernel executes.
@@
Attention Kernel Optimization. Fused, on-chip attention kernel implementations like FlashAttention and FlashInfer reduce global memory traffic by combining attention steps into a single kernel (dao2022flashattention; shah2024flashattention3; ye2025flashinfer), but their query-centric (one-query-per-CTA) design cannot exploit workload-level prefix sharing. Subsequent works (zhu2024relayattention; ye2024chunkattention; lin2024parrot; zheng2024batchllm) pack the single shared system prompt to reduce redundant memory accesses, but can not generalize to workloads with multi-level prefixes. Several recent works (zihao2024cascade; yao2024deft; juravsky2024hydragen) extend to multi-level prefixes but rely on simple packing strategies that overlook trade-offs between overhead and savings. (pan2025fasttree) also addresses multi-level prefix optimization, but its compute-oriented packing cost model mismatches the memory-bound nature of decode attention, leading to suboptimal results. Furthermore, all these works use one-size-fits-all kernel design, leading to resource inefficiency. In contrast, PAT uses a memory-centric packing strategy with multi-tile kernels and multi-stream execution to reduce redundant memory accesses and improve resource utilization for dynamic workloads.
@@
GPU Scheduling. GPU scheduling strategies further improve kernel efficiency by balancing compute and memory demands: (kamath2025pod) fuses prefill and decode via virtual CTAs to expose resource complementarity, and (zhu2025nanoflow) groups SMs so different SMs specialize in specific tasks, raising utilization. These techniques are largely orthogonal to PAT and could be combined to further reduce execution inefficiencies. Work like (song2024tackling) shows that CTA-style scheduling on alternative accelerators (e.g., NPUs) can help eliminate the remaining resource bubbles in PAT ’s execution (§9). Here, we focus on GPU decode-attention optimizations and leave heterogeneous-scheduler integration to future work.
@@ 11. Conclusion
In this work, we present PAT, a prefix-aware attention kernel implementation for LLM decoding. PAT reduces redundant global memory access and improves bandwidth utilization through a pack–forward–merge paradigm. The design incorporates a heuristic pack scheduler, a resource-efficient multi-tile kernel, and forward-stage optimizations that mitigate execution bubbles. We implement PAT as an off-the-shelf plugin to vLLM and demonstrate that it reduces the attention latency by 53.5% on average under synthetic workloads and reduces the TPOT by up to 93.1% under real-world workloads compared with existing works.
@@ 12. Acknowledgments
Appendix A Artifact Appendix
A.1. Abstract
We provide the source code of PAT along with scripts to reproduce experimental results presented in the evaluation section. This appendix includes instructions for reproducing the kernel performance evaluation results in Figure 10 and four representative end-to-end performance evaluation results in Figure 11.
@@ A.1. Abstract
A.2. Artifact check-list (meta-information)
@@
  • How much disk space required (approximately)?: 100GB.
@@
  • How much time is needed to complete experiments (approximately)?: $\approx$2 hours for kernel performance experiment (Figure 10); $\approx$4 hours for the four selected end-to-end performance experiments (Figure 11).
@@
  • Archived (provide DOI)?: 10.5281/zenodo.17943362.
A.3. Description
A.3.1. How to access
A.3.2. Hardware dependencies
@@ A.5. Experiment workflow
    Run the end-to-end serving performance experiments. Note that completing all experiments requires over 60 GPU-hours, so we provide two scripts for convenience: (1) run_e2e_bench_part.sh: runs a subset of experiments (QPS=7, all workloads, all baselines) for quick verification; (2) run_e2e_bench_full.sh: runs all experiments in Figure 11.
@@
    \$ # Quick verification (4-5 GPU-hours)
@@
A.6. Evaluation and expected results
  1. (1)
@@ A.6. Evaluation and expected results
  2. (2)
    Plots similar to Figure 10 and Figure 11 can be found in the fig directory.

### v3

@@ 8.7. Overhead Analysis
Figure 16. Overhead of pack scheduler and pre-attention task latency of serving system.
The primary overhead introduced by PAT comes from the pack scheduler, which packs decode batches into CTAs at runtime. PAT’s lazy update mechanism (§5.1) is designed to mitigate the overhead. It reduces the scheduler’s triggering frequency without affecting correctness and overlaps its execution with the serving system’s pre-attention tasks (e.g., metadata preparation, QKV projection). To validate, we measure both the scheduling latency and the pre-attention task latency under toolagent and conversation traces at request rates of 5 and 8 req/s. As shown in Figure 16, the average scheduling latency is lower than the pre-attention task latency by 42.3% and 49.6%, respectively. Consistently, the P99 TPOT in Figure 12 c1- c4 shows 19.4–93.4% reduction compared with the baselines. Therefore, when running on an asynchronous CPU thread, the pack scheduler will not introduce additional end-to-end latency, demonstrating the effectiveness of the lazy update mechanism in PAT.
@@ 9. Discussion
Prospects and Limitations. PAT leverages cross-request prefix sharing to improve bandwidth utilization and cut redundant global memory accesses in memory-bound decode attention. Its effectiveness depends on three factors: (1) hardware compute-to-bandwidth ratio. As GPUs become more compute-dominant (e.g., NVIDIA V100 to B200: 139 to 312 FLOP/Byte), the gap between compute and memory widens, so memory-focused designs like PAT become increasingly valuable; (2) model architecture. PAT yields large gains for common architectures that retain a KV cache (MHA, GQA), but benefits may shrink for architectures that compress or remove KV state (MLA (Liu et al., 2024), linear attention (Shen et al., 2021), MLKV (Zuhri et al., 2024)); and (3) prefix-sharing ratio in the batch. High concurrency with cross-request shared prefixes amplifies PAT ’s advantage, whereas small batches or workloads without shared prefixes limit the improvement (see Figure 11 1, 1920).
@@ 10. Related Work
KV Cache Related Optimization. Prior work reduced KV cache memory and fragmentation: FasterTransformer (NVIDIA, 2025d) provided an early static-batching implementation, Orca (Yu et al., 2022) improved utilization with continuous batching but relied on pre-allocated caches, which leads to memory fragmentation. vLLM introduced a paged KV cache to cut waste by virtual memory management (Kwon et al., 2023). SGLang’s prefix-reuse lets shared prefix KV blocks be reused across requests and is widely used in production (Zheng et al., 2024b). Recent systems (Qin et al., 2024; Lee et al., 2024; Hu et al., 2025) further extend KV capacity by offloading KV blocks between GPU and CPU or NVMe storage, which reduces on-GPU memory pressure at the cost of additional data movement. These approaches shrink KV memory costs but do not accelerate decode attention, which dominates latency as context and output lengths grow. PAT is orthogonal to these approaches because its pack scheduler only relies on logical KV block IDs, and the serving system transfers the required KV blocks to GPU memory before the attention kernel executes.
@@
Attention Kernel Optimization. Fused, on-chip attention kernel implementations like FlashAttention and FlashInfer reduce global memory traffic by combining attention steps into a single kernel (Dao et al., 2022; Shah et al., 2024; Ye et al., 2025), but their query-centric (one-query-per-CTA) design cannot exploit workload-level prefix sharing. Subsequent works (Zhu et al., 2024; Ye et al., 2024b; Lin et al., 2024a; Zheng et al., 2024a) pack the single shared system prompt to reduce redundant memory accesses, but can not generalize to workloads with multi-level prefixes. Several recent works (Ye et al., 2024a; Yao et al., 2024; Juravsky et al., 2024) extend to multi-level prefixes but rely on simple packing strategies that overlook trade-offs between overhead and savings. (Pan et al., 2025) also addresses multi-level prefix optimization, but its compute-oriented packing cost model mismatches the memory-bound nature of decode attention, leading to suboptimal results. Furthermore, all these works use one-size-fits-all kernel design, leading to resource inefficiency. In contrast, PAT uses a memory-centric packing strategy with multi-tile kernels and multi-stream execution to reduce redundant memory accesses and improve resource utilization for dynamic workloads.
@@
GPU Scheduling. GPU scheduling strategies further improve kernel efficiency by balancing compute and memory demands: (Kamath et al., 2025) fuses prefill and decode via virtual CTAs to expose resource complementarity, and (Zhu et al., 2025) groups SMs so different SMs specialize in specific tasks, raising utilization. These techniques are largely orthogonal to PAT and could be combined to further reduce execution inefficiencies. Work like (Song et al., 2024) shows that CTA-style scheduling on alternative accelerators (e.g., NPUs) can help eliminate the remaining resource bubbles in PAT ’s execution (§9). Here, we focus on GPU decode-attention optimizations and leave heterogeneous-scheduler integration to future work.
@@ 11. Conclusion
In this work, we present PAT, a prefix-aware attention kernel implementation for LLM decoding. PAT reduces redundant global memory accesses and improves bandwidth utilization through a pack–forward–merge paradigm. The design incorporates a heuristic pack scheduler, a resource-efficient multi-tile kernel, and forward-stage optimizations that mitigate execution bubbles. We implement PAT as an off-the-shelf plugin to vLLM and demonstrate that it reduces the attention latency by 53.5% on average under synthetic workloads and reduces the TPOT by up to 93.1% under real-world workloads compared with existing works.
@@ 12. Acknowledgments
Figure 17. Normalized kernel performance (higher is better) of PAT and the baselines for the attention computation across various decode batch configurations on NVIDIA H100 GPU. The four panels from top to bottom show head configurations (num_attention_heads/num_key_value_headsnum\_attention\_heads/num\_key\_value\_heads) of 32/32, 16/8, 32/8, and 64/8. Missing bars arise because RelayAttention lacks support for multi-level or multiple first-level prefixes, and FastTree does not support the 16/8 and 64/8 head settings.
Appendix A Kernel Performance on H100 GPU
PAT’s multi-tile kernel adapts to different GPU architectures by re-deriving equivalent tile-size configurations from the three constraints in §5.2. This allows the kernel to match different memory hierarchies and bandwidths. To validate this, we compare PAT with the baselines in §8.2 on an NVIDIA H100 GPU, using configurations in Figure 11. As in Figure 17, PAT achieves kernel performance of $1.3-6.9\times$ to baselines with shared prefixes ( 1- 18), and achieves consistent performance without shared prefixes ( 19 and 20). These results further confirm the robustness of PAT’s design.
Appendix B Artifact Appendix
B.1. Abstract
We provide the source code of PAT along with scripts to reproduce experimental results presented in §8. This appendix includes instructions for reproducing the kernel performance evaluation results in Figure 11 and four representative end-to-end performance evaluation results in Figure 12.
@@ B.1. Abstract
B.2. Artifact check-list (meta-information)
@@
  • How much disk space required (approximately)?: 200GB.
@@
  • How much time is needed to complete experiments (approximately)?: $\approx$2 hours for kernel performance experiment (Figure 11); $\approx$10 hours for the selected end-to-end performance experiments (Figure 12).
@@
  • Archived (provide DOI)?: 10.5281/zenodo.18217189.
B.3. Description
B.3.1. How to access
B.3.2. Hardware dependencies
@@ B.5. Experiment workflow
    Run the end-to-end serving performance experiments. Note that completing all experiments requires over 60 GPU-hours, so we provide two scripts for convenience: (1) run_e2e_bench_part.sh: runs a subset of experiments (QPS=7&9, all workloads, all baselines) for quick verification; (2) run_e2e_bench_full.sh: runs all experiments in Figure 12.
@@
    \$ # Quick verification (8-10 GPU-hours)
@@
B.6. Evaluation and expected results